Circular No. 7925

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

4U 1543–47 R. W. Hunstead and J. Webb, University of Sydney, report the detec-

tion of a radio counterpart of the x-ray transient 4U 1543–47 (IAUC 7920). A 12-hr integration with the Molonglo Observatory Synthesis Telescope on June 18.50 UT (mid-exposure) yielded a preliminary 843-MHz flux density of 11 ± 1.5 mJy, coincident with the optical position. No source was seen at this position in an archival 1998 observation with a similar noise level.

NEW TRANSITS FROM THE OGLE-III SURVEY A. Udalski, Warsaw University Observatory, on behalf of the OGLE-III

survey team, writes that photometric data collected during the 2001 OGLE planetary/low-luminosity object transit campaign (Udalski et al. 2002, Acta Astron. 52, 1), whereby smaller objects are thought to be transiting across stars, have been reanalyzed for transits with the BLS method (Kovacs et

al. 2002, http://xxx.lanl.gov/abs/astro-ph/?0206099). Thirteen ad-

ditional transit objects were detected, increasing the total number found to 59. The list of new transit objects is available from the OGLE Internet

archive (http://www.astrouw.edu.pl/~ogle; http://bulge.princeton

.edu/~ogle). The shapes of the transit light curves indicate a few new Jupiter-sized objects (R < 1.6R_{J}). The most intriguing case is OGLE-TR-56 (α = 17^{h}56^{m}35^{s}.51, δ = −29^{o}32′21′′.2, equinox 2000.0; I = 15.3, V − I = +1.3). Its transit has a depth of only 0.013 mag and a short orbital period of 1.21192 days (based on ten individual transits). A model transit fit to the data indicates the size of the transiting object to be 0.7R_{J} if the passage were central (i = 90^{o}); i.e., it may be less than Saturn-sized. If confirmed spectroscopically to be a planet, it would have the shortest period and would be one of the smallest extrasolar planets (a ‘hot’ Saturn). Spectroscopic follow-up observations are encouraged. Just two radial-velocity measurements with an accuracy of ∼ 0.5 km/s, taken at phases near 0.25 and 0.75, should place a transiting object in the planetary mass range, if no variation is found. One of the most promising planetary transit candidates discovered during the 2001 campaign, OGLE-TR-40, was monitored with the 1.3-m OGLE telescope at Las Campanas Observatory for the expected transit, and a 0.025-mag drop in brightness (corresponding to the transit) was detected on 2002 June 15.335 UT, ∼ 3 hr earlier than predicted from 2001 data. Timing of the transit allows a refinement of the ephemeris, based now on more than 100 cycles: HJD_{min} = 2452060.03543 + 3.43077E.

2002 June 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green